George and Linda each have 20 dollars in the bank. Each week George adds 3 dollars to his account. Linda adds 1 dollar to her account each week. After how many weeks does George have twice as much as Linda?1
Answer by Theo(13342)

looks like 20 weeks.

in 20 weeks, george will have 20 + 3 * 20 = 20 + 60 = 80
in 20 weeks, linda will have 20 + 1 * 20 = 20 + 20 = 40

80 is 2 times 40.

the formula for george is y = 20 + 3x
the formula for linda is y = 20 + x

x represents the number of weeks.

when george has 2 times as much as linda, then 20 + 3x is equal to 2 * (20 + x)

solve that equation for your answer.

20 + 3x = 2 * (20 + x) becomes:

20 + 3x = 40 + 2x.

subtract 2x from both sides of the equation and subtract 20 from both sides of the equation to get:

3x - 2x = 40 - 20

simplify to get x = 20

that's your solution.